OpenStudy (anonymous):

Which ratio correctly compares 60 mm to 45 cm?

OpenStudy (anonymous):

45 cm is 450 mm. Therefore, 60mm / 45 cm = 60 mm/450 mm = 60/ 450 = 2/15

OpenStudy (anonymous):

60/450 which you can simplify